# Model Generalization on Text Attribute Graphs: Principles with Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have recently been introduced to graph learning, aiming to extend their zero-shot generalization success to tasks where labeled graph data is scarce. Among these applications, inference over text-attributed graphs (TAGs) presents unique challenges: existing methods struggle with LLMs’ limited context length for processing large node neighborhoods and the misalignment between node embeddings and the LLM token space. To address these issues, we establish two key principles for ensuring generalization and derive the framework LLM-BP accordingly: (1) Unifying the attribute space with task-adaptive embeddings, where we leverage LLM-based encoders and task-aware prompting to enhance generalization of the text attribute embeddings; (2) Developing a generalizable graph information aggregation mechanism, for which we adopt belief propagation with LLM-estimated parameters that adapt across graphs. Evaluations on 11 real-world TAG benchmarks demonstrate that LLM-BP significantly outperforms existing approaches, achieving 8.10% improvement with task-conditional embeddings and an additional 1.71% gain from adaptive aggregation. Task-adaptive embeddings and codes are publicly available at [^1] https://github.com/Graph-COM/LLM_BP,
https://huggingface.co/datasets/Graph-COM/Text-Attributed-Graphs
@@ 1 Introduction
Inspired by the remarkable generalization capabilities of foundation models for text and image data [1, 2, 3], researchers have recently explored extending these successes to graph data [4, 5, 6, 7, 8], aiming to develop models that generalize to new or unseen graphs and thereby reduce reliance on costly human annotation [9, 10, 11, 12]. Among various types of graph data, text-attributed graphs (TAGs) have found a wide range of applications. These graphs combine both topological relationships and textual attributes associated with each node, which naturally arises in recommendation systems (where user and item nodes may have textual descriptions or reviews) [13], academic graphs (where publications include extensive textual metadata) [14, 15], and financial networks (where transactions and accounts come with textual records) [16, 17]. Given the labeling challenges posed by cold-start problems in recommendation systems or fraud detection in financial networks, methods that can operate with limited labeled data are crucial. In particular, robust zero-shot node labeling across unseen TAGs has become an area of great interest.
Numerous studies have been dedicated to inference tasks on TAGs. Early efforts have primarily focused on adapting pre-trained language model (LM) encoders [18, 19], sometimes in combination with graph neural networks (GNNs) [20, 21], to incorporate structural information. However, these approaches often struggle to achieve strong generalization performance, largely due to the limited capacity of the underlying models. With the advent of large language models (LLMs) [22, 23], researchers have proposed two main strategies for integrating LLMs into TAG inference: 1) Direct Node-Text Input. Here, raw node texts are directly fed into LLMs. This method demonstrates reasonably good zero-shot performance on TAGs when text attributes are highly informative for node labels [24, 25]. However, when the textual attributes are insufficient, it becomes necessary to aggregate information from a larger neighborhood in the graph, while this is constrained by the limited context length LLMs can digest and reason over. 2) Embedding-Based Integration. In this approach, node texts and their neighboring structural information are first encoded into compressed embeddings, which are then processed by LLMs [26, 27, 28, 29, 30]. Because LLMs are not inherently trained on arbitrary embedding spaces, aligning these embeddings with the LLM’s token space is essential - an idea partly inspired by how vision-language models align multimodal data [3, 31]. However, unlike the vision-language domain, where large-scale text–image pairs [32] are abundant, the graph domain typically lacks comparable datasets. This scarcity reduces the model’s generalization in practice.
@@
In contrast to prior heuristic approaches, this work aims to design a method from first principles for robust zero-shot generalization on TAGs. Because TAGs are inherently multimodal, the proposed method must simultaneously address potential distribution shifts in both textual attributes and graph structure. Specifically, text attributes can vary widely, for example from scientific papers [14] to e-commerce product reviews [33]. Edge connection patterns can range from homophilic graphs such as citation networks, where papers on similar themes are linked [15], to heterophilic graphs such as webpages, which connect nodes with distinct topics [34]. Moreover, the labeling task itself can shift which requires a task-adaptive approach to process both textual features and network structure. Consequently, the core insight behind our model design is grounded in the following two key principles.
@@
Principle I: Unifying the text space and obtaining task-adaptive embeddings. LLMs offer powerful text-understanding capabilities that naturally unify the textual feature space. However, to handle the large-scale graph aggregation discussed later, we require these capabilities to extend beyond raw text to an embedding space. Hence, we propose to adopt LLM-based encoder models such as LLM2Vec [35, 36] for text embedding. Although this approach might appear to be a naive extension of smaller LM-based embedding methods (e.g., those relying on SBERT [37] or RoBERTa [38]), we argue that leveraging the decoder-induced encoder structure of LLMs is essential for achieving task-adaptive embeddings. In particular, we introduce a novel prompting strategy that encodes text attributes conditioned on inference-task descriptions, enabling significantly improved zero-shot inference - an ability not readily achieved by smaller LM-based embeddings.
@@
Principle II: Developing a generalizable adaptive graph information aggregation mechanism. Graph structure determines the node neighboring relationships and thus the information aggregation from which nodes may benefit the inference. Inspired by the belief propagation (BP) algorithm [39] that gives the optimal statistical inference over pairwise correlated random variables, we propose to regard the graph as a Markov Random Field (MRF), each node as a random variable, and mimic BP to aggregate information for node label inference. Because BP is rooted in basic mathematical principles, this approach is widely generalizable. Algorithmic adaptivity across different TAGs hinges on estimating the coupling coefficients in the graphs, which can be done by having LLMs analyze the attributes of sampled pairs of connected nodes. Moreover, this BP-inspired approach naturally adapts to varying levels of text attribute quality: nodes with higher-quality text attributes present greater influence on their neighbors, and vice versa.
@@
By applying the two principles outlined above, we propose our new strategy, LLM-BP, for zero-shot inference over TAGs. LLM-BP does not require any training or fine-tuning. We evaluate LLM-BP on 11 real-world TAG datasets from various domains, including citation networks [14, 15, 40], e-commerce [33], knowledge graphs [34], and webpage networks [41], covering both homophilic and heterophilic graph structures.
@@ 2 Related Works
Tuning Smaller LM Encoders. These methods typically rely on a source-domain graph for training. Notable works include ZeroG [18] that tunes SBert [37] on source datasets to align class-description embeddings with node text, thereby enhancing zero-shot performance on target datasets. Another approach, UniGLM [19], fine-tunes BERT [42] using contrastive learning on source datasets to yield a more generalizable encoder. GNNs trained with UniGLM embeddings in a supervised manner outperform models that directly adopt LM embeddings.
@@
Training GNNs for Generalization. These methods focus on leveraging graph structure in a generalizable manner. Among them, graph self-supervised learning [43] is particularly common for producing representations without labeled data, often employing contrastive learning or masked modeling [21, 20, 44]. GraphMOE is a more recent technique inspired by the success of mixture-of-experts [45], pre-training parallel graph experts targeting different structures or domains [46, 47, 48, 49]. Others also consider LM-GNN co-training including [50, 51] that also follow a constrastive learning idea. Note that, however, all these methods still require training.
@@
In contrast to the above effort that adopts smaller LM encoders, works that involve the use of LLMs are reviewed in the following and may achieve better generalization. More related works including LLM-based data augmentations for GNN training for generalization and LLMs for other graph reasoning tasks can be found in Appendix. 6.
@@
LLMs with Node-Text Input. LLMs being directly fed with raw node texts demonstrates strong zero-shot ability on TAGs [24, 52, 25]. However, they suffer from the limitation of not being able to incorporate graph structural information.
@@
LLMs with Graph-Embedding Input. With smaller LM-encoded node embeddings, various strategies integrate graph structure by aggregating these embeddings, such as neighborhood-tree traversal or concatenating the averaged embeddings from different hops [26, 27, 28], or via pre-trained GNNs [30, 29]. As mentioned earlier, these methods rely on aligning embeddings with the LLMs’ token space. For instance, LLaGA [26] trains a simple MLP on citation networks and [29] employs a linear projector on the ogbn-Arxiv [53] dataset, both using the next-token prediction loss, while [27] adopts self-supervised structure-aware graph matching as the training objective. However, due to limited TAG-domain data, the space alignment in these methods often remains undertrained, leading to degraded performance.
@@
Multi-Task Graph Foundation Models. More ambitious studies aim to generalize across various graph-related tasks within a single framework. Notable approaches include graph prompting [54], which introduces “prompting nodes” to transform diverse graph tasks into a unified format. These frameworks then train GNNs to address the tasks [18, 55, 47] or further integrate LLMs [56, 57]. Although these works are impressive, they still fail to achieve zero-shot performance comparable to those methods that focus on specific graph data domains.
@@ 3.1 Notations and Problem Formulation
The objective is to infer the labels of nodes in TAGs based on the node attributes and graph structure. This study primarily focuses on the zero-shot setting, where no labeled data are assumed to be available in advance. Additionally, a few-shot setting is considered, where $k$ labeled nodes are known for each class. Due to space limitations, results for the few-shot setting are provided in Appendix 9.4.
@@ 3.2 Motivation and the Overall Framework
LLMs are commonly used as decoders for next-token prediction. While LLMs excel at processing natural language inputs, they are not inherently compatible with graph data. Recently, some studies have explored methods to integrate graph data into LLMs, primarily for reasoning tasks [58, 59, 60].
In the context of TAGs, accurate node label inference relies on effectively combining the attributes of multiple nodes, especially when a node’s individual attributes are insufficient to determine its label. However, as noted earlier, LLMs are constrained by limited context windows, making it challenging to process all attributes from the potentially large set of connected nodes. Traditional approaches to compressing graph structural information involve creating embeddings, such as using GNNs to aggregate information from the target node’s neighbors. While effective, these embedding methods do not seamlessly integrate with LLM inputs and often require non-trivial training effort to align the LLMs’ token space with the node embedding space [26, 29, 27].
Our approach, LLM-BP, does not confine LLMs to their traditional usage. We first leverage their capabilities to generate task-adaptive node embeddings. Then, instead of requiring LLMs to directly process these embeddings, LLMs are further employed to analyze graph data and provide generalizable guidance in aggregating these embeddings. These two steps are to match the two generalization principles proposed in Sec. 1. Classification is ultimately performed by computing the cosine similarity between the final node embeddings $\mathbf{h}^{X}=[h_{1}^{X},...,h_{n}^{X}]$ and candidate class embeddings $\mathbf{q}^{C}=[q_{1}^{C},...,q_{c}^{C}]$. In the zero-shot setting, class embeddings are generated as follows: we randomly sample $l\ll n$ nodes and employ LLMs to infer their labels. The embeddings of sampled nodes form distinct clusters based on LLMs’ prediction. We compute the average embedding of the embeddings closest to the cluster center to obtain the class embedding. In the few-shot setting, class embeddings are obtained by averaging the embeddings of labeled nodes within each class. See Appendix. 7.2 for details.
@@ 3.3 Principle I: Task-Adaptive Node Embeddings $\mathbf{h}^{X}$
Creating generalizable text embeddings is no longer a significant challenge. Even smaller LM encoders, such as SBert [37], are capable of achieving this. Indeed, most existing works utilize these encoders to generate initial node embeddings for TAGs [10, 26, 27, 29]. However, for these embeddings to be directly usable for label prediction without the need for additional transformation models, it is crucial to incorporate task-specific information. In other words, the embeddings must be tailored to the specific task, resulting in what we term task-adaptive embeddings.
Achieving task adaptivity, however, presents a notable challenge. Smaller LM encoders lack the expressive power necessary to encode nuanced task-specific information. This limitation motivates our adoption of LLM-induced encoders, driven by the emergent capabilities of LLMs in contextual learning [61, 62].
There have been recent advancements in extending LLMs to generate text embeddings [35, 63]. In our approach, we utilize a form of LLM2Vec [35], which transforms LLM decoders into encoders via retaining the unidirectional attention mechanism. Following the methodology in [36], we extract the output embedding of $\langle\text{response}\rangle$ the token positioned immediately after the inputs - as the text embedding for the input node attributes.

### v2

# Generalization Principles for Inference over Text-Attributed Graphs with Large Language Models

@@ Abstract
Large language models (LLMs) have recently been introduced to graph learning, aiming to extend their zero-shot generalization success to tasks where labeled graph data is scarce. Among these applications, inference over text-attributed graphs (TAGs) presents unique challenges: existing methods struggle with LLMs’ limited context length for processing large node neighborhoods and the misalignment between node embeddings and the LLM token space. To address these issues, we establish two key principles for ensuring generalization and derive the framework LLM-BP accordingly: (1) Unifying the attribute space with task-adaptive embeddings, where we leverage LLM-based encoders and task-aware prompting to enhance generalization of the text attribute embeddings; (2) Developing a generalizable graph information aggregation mechanism, for which we adopt belief propagation with LLM-estimated parameters that adapt across graphs. Evaluations on 11 real-world TAG benchmarks demonstrate that LLM-BP significantly outperforms existing approaches, achieving 8.10% improvement with task-conditional embeddings and an additional 1.71% gain from adaptive aggregation. The code [^2] https://github.com/Graph-COM/LLM_BP and task-adaptive embeddings [^3] https://huggingface.co/datasets/Graph-COM/Text-Attributed-Graphs are publicly available.
Machine Learning, ICML
@@ 1 Introduction
Inspired by the remarkable generalization capabilities of foundation models for text and image data (Achiam et al., 2023; Liu et al., 2021; Radford et al., 2021), researchers have recently explored extending these successes to graph data (Liu et al., 2023b; Mao et al., 2024; Zhao et al., 2023a; Fan et al., 2024; He et al., 2024), aiming to develop models that generalize to new or unseen graphs and thereby reduce reliance on costly human annotation (Li et al., 2024f; Chen et al., 2024d; Feng et al., 2024; Li et al., 2024g). Among various types of graph data, text-attributed graphs (TAGs) have found a wide range of applications. These graphs combine both topological relationships and textual attributes associated with each node, which naturally arises in recommendation systems (where user and item nodes may have textual descriptions or reviews) (Bobadilla et al., 2013), academic graphs (where publications include extensive textual metadata) (McCallum et al., 2000; Giles et al., 1998), and financial networks (where transactions and accounts come with textual records) (Kumar et al., 2016, 2018). Given the labeling challenges posed by cold-start problems in recommendation systems or fraud detection in financial networks, methods that can operate with limited labeled data are crucial. In particular, robust zero-shot node labeling across unseen TAGs has become an area of great interest.
@@
Numerous studies have been dedicated to inference tasks on TAGs. Early efforts have primarily focused on adapting pre-trained language model (LM) encoders (Li et al., 2024e; Fang et al., 2024), sometimes in combination with graph neural networks (GNNs) (Hou et al., 2022; Veličković et al., 2018), to incorporate structural information. However, these approaches often struggle to achieve strong generalization performance, largely due to the limited capacity of the underlying models. With the advent of large language models (LLMs) (Kaplan et al., 2020; Huang & Chang, 2022), researchers have proposed two main strategies for integrating LLMs into TAG inference: 1) Direct Node-Text Input. Here, raw node texts are directly fed into LLMs. This method demonstrates reasonably good zero-shot performance on TAGs when text attributes are highly informative for node labels (Chen et al., 2024c; Li et al., 2024d). However, when the textual attributes are insufficient, it becomes necessary to aggregate information from a larger neighborhood in the graph, while this is constrained by the limited context length LLMs can digest and reason over. 2) Embedding-Based Integration. In this approach, node texts and their neighboring structural information are first encoded into compressed embeddings, which are then processed by LLMs (Chen et al., 2024b; Tang et al., 2024a; Luo et al., 2024; Wang et al., 2024; Zhang et al., 2024a). Because LLMs are not inherently trained on arbitrary embedding spaces, aligning these embeddings with the LLM’s token space is essential - an idea partly inspired by how vision-language models align multimodal data (Radford et al., 2021; Zhai et al., 2023). However, unlike the vision-language domain, where large-scale text–image pairs (Schuhmann et al., 2022) are abundant, the graph domain typically lacks comparable datasets. This scarcity reduces the model’s generalization in practice.
@@
In contrast to prior heuristic approaches, this work aims to design a method from first principles for robust zero-shot generalization on TAGs. Because TAGs are inherently multimodal, the proposed method must simultaneously address potential distribution shifts in both textual attributes and graph structure. Specifically, text attributes can vary widely, for example from scientific papers (McCallum et al., 2000) to e-commerce product reviews (Ni et al., 2019). Edge connection patterns can range from homophilic graphs such as citation networks, where papers on similar themes are linked (Giles et al., 1998), to heterophilic graphs such as webpages, which connect nodes with distinct topics (Mernyei & Cangea, 2020). Moreover, the labeling task itself can shift which requires a task-adaptive approach to process both textual features and network structure. Consequently, the core insight behind our model design is grounded in the following two key principles.
@@
Principle I: Unifying the text space and obtaining task-adaptive embeddings. LLMs offer powerful text-understanding capabilities that naturally unify the textual feature space, and have recently been shown to benefit from task-adaptive prompting (Kong et al., 2024). However, to handle the large-scale graph aggregation discussed later, we require these capabilities to extend beyond raw text to an embedding space. Hence, we propose to adopt LLM-based encoder models such as LLM2Vec (BehnamGhader et al., 2024; Li et al., 2024a) for text embedding. Although this approach might appear to be a naive extension of smaller LM-based embedding methods (e.g., those relying on SBERT (Reimers, 2019) or RoBERTa (Liu, 2019)), we argue that leveraging the decoder-induced encoder structure of LLMs is essential for achieving task-adaptive embeddings. In particular, we introduce a novel prompting strategy that encodes text attributes conditioned on inference-task descriptions, enabling significantly improved zero-shot inference - an ability not readily achieved by smaller LM-based embeddings.
@@
Principle II: Developing a generalizable adaptive graph information aggregation mechanism. Graph structure determines the node neighboring relationships and thus the information aggregation from which nodes may benefit the inference. Inspired by the belief propagation (BP) algorithm (Murphy et al., 2013) that gives the optimal statistical inference over pairwise correlated random variables, we propose to regard the graph as a Markov Random Field (MRF), each node as a random variable, and mimic BP to aggregate information for node label inference. Because BP is rooted in basic mathematical principles, this approach is widely generalizable. Algorithmic adaptivity across different TAGs hinges on estimating the coupling coefficients in the graphs, which can be done by having LLMs analyze the attributes of sampled pairs of connected nodes. Moreover, this BP-inspired approach naturally adapts to varying levels of text attribute quality: nodes with higher-quality text attributes present greater influence on their neighbors, and vice versa.
@@
By applying the two principles outlined above, we propose our new strategy, LLM-BP, for zero-shot inference over TAGs. LLM-BP does not require any training or fine-tuning. We evaluate LLM-BP on 11 real-world TAG datasets from various domains, including citation networks (McCallum et al., 2000; Giles et al., 1998; Sen et al., 2008), e-commerce (Ni et al., 2019), knowledge graphs (Mernyei & Cangea, 2020), and webpage networks (Craven et al., 1998), covering both homophilic and heterophilic graph structures.
@@ 2 Related Works
Tuning Smaller LM Encoders. These methods typically rely on a source-domain graph for training. Notable works include ZeroG (Li et al., 2024e) that tunes SBert (Reimers, 2019) on source datasets to align class-description embeddings with node text, thereby enhancing zero-shot performance on target datasets. Another approach, UniGLM (Fang et al., 2024), fine-tunes BERT (Kenton & Toutanova, 2019) using contrastive learning on source datasets to yield a more generalizable encoder. GNNs trained with UniGLM embeddings in a supervised manner outperform models that directly adopt LM embeddings.
@@
Training GNNs for Generalization. These methods focus on leveraging graph structure in a generalizable manner. Among them, graph self-supervised learning (Liu et al., 2022) is particularly common for producing representations without labeled data, often employing contrastive learning or masked modeling (Veličković et al., 2018; Hou et al., 2022; Zhao et al., 2024). GraphMOE is a more recent technique inspired by the success of mixture-of-experts (Shazeer et al., 2017), pre-training parallel graph experts targeting different structures or domains (Hou et al., 2024; Liu et al., 2024; Xia & Huang, 2024; Qin et al., 2023). Others also consider LM-GNN co-training including (He & Hooi, 2024; Zhu et al., 2024) that also follow a constrastive learning idea. Note that, however, all these methods still require training.
@@
In contrast to the above effort that adopts smaller LM encoders, works that involve the use of LLMs are reviewed in the following and may achieve better generalization. More related works including LLM-based data augmentations for GNN training for generalization and LLMs for other graph reasoning tasks can be found in Appendix. A.
@@
LLMs with Node-Text Input. LLMs being directly fed with raw node texts demonstrates strong zero-shot ability on TAGs (Chen et al., 2024c; Huang et al., 2024; Li et al., 2024d). However, they suffer from the limitation of not being able to incorporate graph structural information.
@@
LLMs with Graph-Embedding Input. With smaller LM-encoded node embeddings, various strategies integrate graph structure by aggregating these embeddings, such as neighborhood-tree traversal or concatenating the averaged embeddings from different hops (Chen et al., 2024b; Tang et al., 2024a; Luo et al., 2024), or via pre-trained GNNs (Zhang et al., 2024a; Wang et al., 2024). As mentioned earlier, these methods rely on aligning embeddings with the LLMs’ token space. For instance, LLaGA (Chen et al., 2024b) trains a simple MLP on citation networks and (Wang et al., 2024) employs a linear projector on the ogbn-Arxiv (Hu et al., 2020) dataset, both using the next-token prediction loss, while (Tang et al., 2024a) adopts self-supervised structure-aware graph matching as the training objective. However, due to limited TAG-domain data, the space alignment in these methods often remains undertrained, leading to degraded performance.
@@
Multi-Task Graph Foundation Models. More ambitious studies aim to generalize across various graph-related tasks within a single framework. Notable approaches include graph prompting (Liu et al., 2023c), which introduces “prompting nodes” to transform diverse graph tasks into a unified format. These frameworks then train GNNs to address the tasks (Li et al., 2024e; Liu et al., 2023a, 2024) or further integrate LLMs (Yang et al., 2024; Kong et al., 2024). Although these works are impressive, they still fail to achieve zero-shot performance comparable to those methods that focus on specific graph data domains.
@@ 3.1 Notations and Problem Formulation
The objective is to infer the labels of nodes in TAGs based on the node attributes and graph structure. This study primarily focuses on the zero-shot setting, where no labeled data are assumed to be available in advance. Additionally, a few-shot setting is considered, where $k$ labeled nodes are known for each class. Due to space limitations, results for the few-shot setting are provided in Appendix D.6.
@@ 3.2 Motivation and the Overall Framework
LLMs are commonly used as decoders for next-token prediction. While LLMs excel at processing natural language inputs, they are not inherently compatible with graph data. Recently, some studies have explored methods to integrate graph data into LLMs, primarily for reasoning tasks (Perozzi et al., 2024; Zhang et al., 2024c; Tang et al., 2024b).
In the context of TAGs, accurate node label inference relies on effectively combining the attributes of multiple nodes, especially when a node’s individual attributes are insufficient to determine its label. However, as noted earlier, LLMs are constrained by limited context windows, making it challenging to process all attributes from the potentially large set of connected nodes. Traditional approaches to compressing graph structural information involve creating embeddings, such as using GNNs to aggregate information from the target node’s neighbors. While effective, these embedding methods do not seamlessly integrate with LLM inputs and often require non-trivial training effort to align the LLMs’ token space with the node embedding space (Chen et al., 2024b; Wang et al., 2024; Tang et al., 2024a).
Our approach, LLM-BP, does not confine LLMs to their traditional usage. We first leverage their capabilities to generate task-adaptive node embeddings. Then, instead of requiring LLMs to directly process these embeddings, LLMs are further employed to analyze graph data and provide generalizable guidance in aggregating these embeddings. These two steps are to match the two generalization principles proposed in Sec. 1. Classification is ultimately performed by computing the cosine similarity between the final node embeddings $\mathbf{h}^{X}=[h_{1}^{X},...,h_{n}^{X}]$ and candidate class embeddings $\mathbf{q}^{C}=[q_{1}^{C},...,q_{c}^{C}]$. In the zero-shot setting, class embeddings are generated as follows: we randomly sample $l\ll n$ nodes and employ LLMs to infer their labels. The embeddings of sampled nodes form distinct clusters based on LLMs’ prediction. We compute the average embedding of the embeddings closest to the cluster center to obtain the class embedding. In the few-shot setting, class embeddings are obtained by averaging the embeddings of labeled nodes within each class. See Appendix. B.2 for details.
@@ 3.3 Principle I: Task-Adaptive Node Embeddings $\mathbf{h}^{X}$
Creating generalizable text embeddings is no longer a significant challenge. Even smaller LM encoders, such as SBert (Reimers, 2019), are capable of achieving this. Indeed, most existing works utilize these encoders to generate initial node embeddings for TAGs (Chen et al., 2024d, b; Tang et al., 2024a; Wang et al., 2024). However, for these embeddings to be directly usable for label prediction without the need for additional transformation models, it is crucial to incorporate task-specific information. In other words, the embeddings must be tailored to the specific task, resulting in what we term task-adaptive embeddings.
Achieving task adaptivity, however, presents a notable challenge. Smaller LM encoders lack the expressive power necessary to encode nuanced task-specific information. This limitation motivates our adoption of LLM-induced encoders, driven by the emergent capabilities of LLMs in contextual learning (Sahoo et al., 2024; Chen et al., 2023).
There have been recent advancements in extending LLMs to generate text embeddings (BehnamGhader et al., 2024; Muennighoff et al., 2022). In our approach, we utilize a form of LLM2Vec (BehnamGhader et al., 2024), which transforms LLM decoders into encoders via retaining the unidirectional attention mechanism. Following the methodology in (Li et al., 2024a), we extract the output embedding of $\langle\text{response}\rangle$ the token positioned immediately after the inputs - as the text embedding for the input node attributes.